Circular No. 7628

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

NOVA AQUILAE 2001 C. R. Benn, M. Vigotti and J. Holt, Instituto de Astrofísica de Ca-

narias, write that a spectrum of the object in Aquila (IAUC 7627), obtained with the 2.5-m Isaac Newton Telescope (IDS spectrograph, resolution 1 nm) on May 18.118 UT, is clearly that of a nova, with strong Hα, Hβ, Hγ and the usual He and Fe lines. The signal-to-noise on the spectrum is ∼ 100 per pixel step in wavelength.

O. Shemmer, Department of Astrophysics, Tev-Aviv University, reports that a spectrum (range 400–780 nm) obtained using the Wise Observatory 1-m telescope (+FOSC spectrograph) on May 18.88 UT shows several characteristic nova features, such as strong Balmer lines with a mean FWHM of 1100 km/s, He I emission lines at 706.5, 667.8 and 587.6 nm; also several strong Fe II emission lines, such as 492.3 and 501.8 nm. The Hα/Hβ flux ratio is 14. A steep red continuum was also noted.

K. Takamizawa, Saku-machi, Nagano prefecture, communicates the following magnitude information from his photographic patrol films: 2000 Oct. 27.42 UT, [15.1; 2001 Feb. 20.79, [13.1; 26.79, 13.2; Mar. 18.72, 13.0; 21.77, 12.7; 31.74, 12.7; Apr. 26.68, 12.8; May 12.59, 10.8; 16.63, 13.0.

M. Uemura and T. Kato, Kyoto University; and H. Yamaoka, Kyushu University, write that unfiltered CCD images obtained with the Kyoto 0.25m telescope on May 17.79 showed the object at mag ∼ 11.8 (GSC system) in the position α = 19^{h}07^{m}28^{s}.42, δ = +11^{o}44′45′′.8 (equinox 2000.0; GSC-ACT, epoch 1983.609, fitting error 0′′.3). A faint star exists on the digital sky survey (both generations, R and B_{j}) within 0′′.5 of the above position. Its magnitude, compared with USNO A2.0 stars, is r ∼ 18.7 and b ∼ 19.6.”

SUPERNOVA 2001bq IN NGC 5534 S. Nakano, Sumoto, Japan, communicates that K. Itagaki, Teppo-

machi, Yamagata (0.60-m f/6.0 reflector + CCD), reports his discovery of a supernova (mag 15.9) in NGC 5534 on five 15-s frames taken on May 17.587 UT and confirmed on May 12.562 (mag 16.0), as well as on May 18.566, when the position was measured as α = 14^{h}17^{m}42^{s}.06, δ = −7^{o}25′01′′.6 (equinox 2000.0). W. D. Li, University of California at Berkeley, reports that SN 2001bq was confirmed on unfiltered images taken with the Katzman Automatic Imaging Telescope on May 10.3 UT (mag ∼ 16.0) and 14.3 (mag ∼ 15.3). Position end figures were measured as 42^{s}.14, 00′′.7, which is 27′′.6 east and 1′′.5 north of the nucleus of NGC 5534. An image taken with the Tenagra Observatory 0.5-m automatic telescope by M. Schwartz on Apr. 29.3 showed nothing at the position (limiting mag ∼ 19.5).

2001 May 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Brian G. Marsden